# Joint Diffusion models in Continual Learning

### v1

@@ 1 Introduction
Contemporary deep neural networks can be trained to human-level performance on a variety of different problems. However, contrary to humans, when retraining the same models to learn an additional task, neural networks suffer from catastrophic forgetting [16] defined as an abrupt loss in performance on the previous task when adapting to a new one. Continual learning approaches try to mitigate this issue. In particular, replay-based techniques retrain the model with a mix of new data examples and samples from the past tasks stored in the memory buffer [47, 1]. However, since the memory buffer has to constantly grow, new methods [57, 65] use generative models to generate synthetic version of past experiences.
@@
Train a generative model $G$ and classifier $C$ using data $D_{1}$
for each subsequent task $i$ do
     Sample dataset $S$ with generations from the generative model $G$
     Form a combined dataset $B\leftarrow D_{i}\cup S$ with new data from task $i$ and the sampled generations
     Update the classifier $C$ with the combined dataset $B$
     Update the generative model $G$ with the combined dataset $B$
@@
Such an approach introduces an imbalance between the generative model and the classifier. The generative model plays a crucial role in a knowledge consolidation method that stores information about old tasks, while the classifier is highly dependent on the quality of generated samples, and the effectiveness of the knowledge transfer from the generative part. In practice, several works highlight that it might even be beneficial to completely reset the weights of the classifier, and train it from scratch at each task, using the mix of new data and sampled generations [11, 71, 32].
@@
In this work, we propose to address this limitation by introducing a joint model that combines the generative and discriminative models into a single parametrization. In particular, we extract the latent features from the UNet [48] trained as a denoising diffusion model and use them for classification. Following [10], we optimize our joint model with a sum of generative and discriminative objectives. Such an approach directly removes the need for transferring the knowledge from the generative model to the classifier when retraining the model on a sequence of tasks.
Figure 1: Joint modeling limits degradation caused by retraining the model on generated samples.
To visualize this issue, we conduct an experiment that highlights the drawback of recent approaches. Using only data from the first task of the CIFAR100 setup, we train the classifier and our joint-diffusion model on the training dataset until convergence (for 110 epochs). Then, we continue to optimize the models, but instead of using real data, we sample data from the diffusion model. As presented in Fig. 1, the performance of the classifier (blue line), drastically decreases as the effect of the rehearsal with generated samples. At the same time, the performance of our joint diffusion (green) converges after a small drop in performance.This indicates that generating rehearsal samples from the same model that we update, facilitates knowledge retention in generative replay.
On top of this observation, we adapt the joint-diffusion models to the continual learning scenario, by combining it with knowledge distillation and two-stage training approaches, and evaluate it in several continual learning tasks. In our experiments, we show that JDCL achieves state-of-the-art performance when compared with other generative replay methods. Moreover, we show that our approach can be easily applied in the continual semi-supervised setup, where thanks to the shared parametrization we can observe an alignment between labeled and unlabeled data. In this task, we show the superiority of our method over the buffer-based techniques. Finally, we also show the quality of representations built by our method by evaluating it in a self-supervised manner.
The main contribution of this work can be summarized as follows:
@@
  • We show that in class-incremental setup, our approach outperforms recent state-of-the-art generative replay techniques and approaches the soft-upper bound defined by a method with an infinite memory buffer.
@@ 2 Related Works
Several works further extend the basic generative replay technique. In [65], followed by [35], authors propose to replay internal data representations instead of the original inputs. This approach improves model stability but limits its adaptation to new tasks. On the other hand, several methods drew inspiration from neuroscience and introduced techniques that mimic human memory system by dividing the update of the model into two [23, 12] or even three stages [68].
Recently, diffusion-based generative models (DDGMs) [59, 20] have come to the forefront as a state-of-the-art generative method. Therefore, there is a growing research on continual learning with DDGMs. Zajac et al. [76] provide an overview of how DDGMs work when combined with baseline CL methods. In [18] authors employ DDGMs with classifier guidance [13] as a source of rehearsal examples. This idea is further extended in [9], where authors overview different sampling strategies with DDGMs. There are several methods that explore the possibilities for efficient updates of large-scale text-to-image diffusion models. In DreamBooth [51] authors introduce a method for finetuning text-to-image DDGMs with Prior Preservation Loss, that employs generative replay to mitigate catastrophic forgetting. Similarly, [58] extend this idea with Low Rank Adaptation (LORA [21]) technique, while in [56] apply this method for object and style update. On the other hand, Basu et al. [3], show that knowledge in text-to-image models is highly localized and can be updated directly in the attention layers of the text encoder.
@@
Semi-supervised methods confront the problem of training models on large datasets where only a few samples have associated annotations [60, 75]. Recently, most of the best-performing semi-supervised methods belong to the family of hybrid methods that combine consistency regularization and pseudo-labeling  [4, 28, 60, 75, 77]. The main idea behind consistency regularization is an assumption that the model should yield similar outputs when fed perturbed versions of the same image or when its weights are slightly perturbed [30, 53, 62, 40]. On the other hand, pseudo-labeling techniques exploit the concept of using synthetic labels for unlabeled data. More precisely, this involves utilizing "hard" labels (i.e., the arg max of the model’s output) for which the predicted class probability surpasses a predetermined threshold [73, 7, 70, 46, 31, 75]. Apart from consistency regularization and pseudo-labeling, a third group of SSL methods employs additional deep generative models  [75] such as Variational Auto-Encoder (VAE) [33, 43, 14, 38, 25], or Generative Adversarial Networks (GAN)  [54, 61, 72, 17, 36], to learn the underlying class-conditional distribution of partially labeled data.
@@ 3.2 Joint Diffusion Model
$L_{\text{class}}(\nu,\omega)=-\mathbb{E}_{\mathbf{x}_{0},y}\left[\sum_{k=0}^{K-1}\mathds{1}[y=k]\log\frac{\exp\left(\varphi_{k}\right)}{\sum_{c=0}^{K-1}\exp\left(\varphi_{c}\right)}\right],$ (6)
@@
$L_{JD}(\nu,\psi,\omega)=\alpha\cdot L_{\text{class}}(\nu,\omega)-L_{0}(\nu,\psi)-\sum_{t=2}^{T}L_{t,\text{diff}}(\nu,\psi)-L_{T}(\nu,\psi),$ (7)
@@ 4 Method
Figure 2: Overview of JDCL. To adapt the global model to task $t+1$, we first train a local joint diffusion model on task $t+1$ only. We then use the global model and the local model to generate a synthetic dataset consisting of samples from tasks $11,2\dots t+$ 1. Finally, we fine-tune the global model with the generated dataset and a combination of joint diffusion and knowledge distillation losses.
@@
To facilitate knowledge consolidation in JDCL, we additionally propose to split the training process into two parts: a local and a global training stage. Our procedure begins with the former, where we train the joint diffusion model $p_{\nu,\psi,\omega}(x_{0:T},y)$ on the currently available dataset $\tau_{t}$. We call that model local and we optimize it with the loss provided in Equation 7.
Following that, we move to the global stage. We begin by generating a rehearsal dataset $S_{1\dots t-1}$ for tasks $1$ to $t-1$ by sampling from the frozen version of the global joint diffusion model. We then combine this dataset with a synthetic dataset $S_{t}$, which we generate with the local model. All the samples are generated unconditionally by the diffusion part of the joint diffusion. The labels for those generations are then assigned by the classifier part of the joint model.
Finally, to acquire the next global model, we fine-tune the previous global joint diffusion on the combined synthetic dataset with a joint diffusion loss $L_{JD}$ defined in Equation 7.
@@
$L_{JDKD}(\nu,\psi,\omega;p^{f})=\alpha_{KD}\cdot L^{KD}_{\text{class}}(\nu,\omega)-L^{KD}_{0}(\nu,\psi)-\sum_{t=2}^{T}L^{KD}_{t,\text{diff}}(\nu,\psi)-L^{KD}_{T}(\nu,\psi),$
@@
Finally, the full training objective, which we optimize to fine-tune the global model for task $t$ can be expressed as:
@@
$L_{\text{ssl}}(\nu,\omega)=-\mathbb{E}_{\mathbf{x}_{0},y}\{\mathds{1}[max(\varphi_{q})\geq\tau]H(\hat{q_{b}},g_{\omega}(e_{\nu}(A(x_{\text{u}}))))\}$ (10)
@@
$L(\nu,\psi,\omega)=\alpha\cdot L_{\text{class}}(\nu,\omega)+\alpha\cdot L_{\text{ssl}}(\nu,\omega)-L_{0}(\nu,\psi)-\sum_{t=2}^{T}L_{t,\text{diff}}(\nu,\psi)-L_{T}(\nu,\psi).$
@@ 5.1 Experimental Setup
We evaluate our method on two commonly used datasets – CIFAR10 and CIFAR100 [27]. Both of those datasets contain 50000 training and 10000 test images of size $32\times 32$ divided in 10 or 100 classes. For class-incremental setup, we split the CIFAR-10 dataset into 5 equal tasks and the CIFAR-100 dataset into 5 or 10 equal tasks. For semi-supervised setups, we train our models on two different ratios of labeled data in the dataset i.e. 0.8% and 5%. Those ratios correspond to 4 and 25 labelled examples per class in CIFAR100, and to 40 and 250 samples for CIFAR10.
@@
We use the same joint diffusion model architecture in all experiments. For the UNet, we follow the implementation provided in [13]. For the classifier, we use a 2-layer feed-forward net with a hidden layer of 1024 and a LeakyRELU activation function. The detailed description of the training details is provided in Appendix A, and in the code repository [^1] Hidden for review.
@@ 5.2 Main results
Incremental setup
Method CIFAR-10 CIFAR-100 CIFAR-10 CIFAR-100
$T=5$ $T=5$ $T=10$ $T=5$ $T=5$ $T=10$
Joint 93.14 $\pm$ 0.16 72.32 $\pm$ 0.24 - - -
Continual Joint 86.41 $\pm$ 0.32 73.07 $\pm$ 0.01 64.15 $\pm$ 0.98 02.90 $\pm$ 0.08 07.80 $\pm$ 0.55 06.67 $\pm$ 0.36
Fine-tuning 18.95 $\pm$ 0.20 16.92 $\pm$ 0.03 9.12 $\pm$ 0.04 94.65 $\pm$ 0.17 80.75 $\pm$ 0.22 87.67 $\pm$ 0.07
DGR VAE 28.23 $\pm$ 3.84 19.66 $\pm$ 0.27 10.04 $\pm$ 0.17 57.21 $\pm$ 9.82 42.10 $\pm$ 1.40 60.31 $\pm$ 4.80
DGR+distill 27.83 $\pm$ 1.20 21.38 $\pm$ 0.61 13.94 $\pm$ 0.13 43.43 $\pm$ 2.60 29.30 $\pm$ 0.40 21.15 $\pm$ 1.30
RTF 30.36 $\pm$ 1.40 17.45 $\pm$ 0.28 12.80 $\pm$ 0.78 51.77 $\pm$ 1.00 47.68 $\pm$ 0.80 45.21 $\pm$ 5.80
MeRGAN 51.65 $\pm$ 0.40 09.65 $\pm$ 0.14 12.34 $\pm$ 0.15 - - -
BIR 36.41 $\pm$ 0.82 21.75 $\pm$ 0.08 15.26 $\pm$ 0.49 65.28 $\pm$ 1.27 48.38 $\pm$ 0.44 53.08 $\pm$ 0.75
GFR 26.70 $\pm$ 1.90 34.80 $\pm$ 0.26 21.90 $\pm$ 0.14 49.29 $\pm$ 6.03 19.16 $\pm$ 0.55 17.44 $\pm$ 2.20
DDGR 43.69 $\pm$ 2.60 28.11 $\pm$ 2.58 15.99 $\pm$ 1.08 62.51 $\pm$ 3.84 60.62 $\pm$ 2.13 74.70 $\pm$ 1.79
DGR diffusion 59.00 $\pm$ 0.57 28.25 $\pm$ 0.22 15.90 $\pm$ 1.01 40.38 $\pm$ 0.32 68.70 $\pm$ 0.65 80.38 $\pm$ 1.34
GUIDE 64.47 $\pm$ 0.45 41.66 $\pm$ 0.40 26.13 $\pm$ 0.29 24.84 $\pm$ 0.05 44.30 $\pm$ 1.10 60.54 $\pm$ 0.82
JDCL 83.69 $\pm$ 1.44 47.95 $\pm$ 0.61 29.04 $\pm$ 0.41 11.13 $\pm$ 3.27 37.20 $\pm$ 0.75 59.55 $\pm$ 0.66
To compare our method with other generative rehearsal approaches, we measure its performance on several CL benchmarks. In Tab. 1, we show that our method outperforms other evaluated generative replay techniques in terms of average accuracy after the last task. In particular, JDCL outperforms the recent state-of-the-art method–GUIDE on both average accuracy and forgetting. On CIFAR10, our method approaches the soft upper limit, calculated as a continual joint method trained with an infinitely long replay buffer. On this benchmark, we surpass the previous state-of-the-art score by over 19 points, which is an increase of around 30%.
@@
Semi-supervised continual learning
Apart from standard supervised setup, we evaluate our method on a set of semi-supervised continual setups and compare it to other approaches. As shown in Tab. 2 JDCL outperforms other methods on all CIFAR10 setups, while also achieving reasonable accuracy on CIFAR100. What’s worth noting is that JDCL is able to reach state-of-the-art performance on CIFAR10 without the need to store any data from the previous tasks.

### v2

@@ 1 Introduction
Contemporary deep neural networks can be trained to human-level performance on a variety of different problems. However, contrary to humans, when retraining the same models to learn an additional task, neural networks suffer from catastrophic forgetting [16] defined as an abrupt loss in performance on the previous task when adapting to a new one. Continual learning approaches try to mitigate this issue. In particular, replay-based techniques retrain the model with a mix of new data examples and samples from the past tasks stored in the memory buffer [48, 1]. However, since the memory buffer has to constantly grow, new methods [58, 68] use generative models to generate synthetic versions of past experiences.
@@
Input: $G$, $C$, sequence of tasks $\{D_{i}\}_{i=1}^{N}$
$G\leftarrow$ train_generative_model($G,D_{1}$)
$C\leftarrow$ train_classifier($C,D_{1}$)
for i in 2 …$N$ do
    $S\leftarrow$ sample_dataset($G$)
    $B\leftarrow D_{i}\cup S$
    $G\leftarrow$ train_generative_model($G$, $B$)
    $C\leftarrow$ train_classifier($C$, $B$)
@@
Such an approach introduces an imbalance between the generative model and the classifier. The generative model plays a crucial role as a knowledge consolidation method that stores information about old tasks, while the classifier is highly dependent on the quality of generated samples, and the effectiveness of the knowledge transfer from the generative part of the system. Consequently, several studies [9, 72, 33] suggest that it may be advantageous to reset the classifier’s weights entirely and retrain it from scratch for each task, using a blend of new data and sampled generations – a solution that contradicts the principles of continual learning.
@@
In this work, we propose to address this limitation by introducing a joint model that combines the generative and discriminative models into a single parametrization. In particular, we extract the latent features from the UNet [49] trained as a denoising diffusion model and use them for classification. We optimize our joint model with a sum of generative and discriminative objectives. Such an approach directly removes the need for transferring the knowledge from the generative model to the classifier when retraining the model on a sequence of tasks.
Figure 1: Further training of a classifier using data generated from the diffusion model trained with the same dataset harms the classifier performance (blue line). However, joint modeling, especially in combination with knowledge distillation significantly limits this degradation and retains almost the initial performance. We report accuracy averaged across three different seeds.
To visualize the main benefit of this approach, we conduct a simple experiment that highlights the drawbacks of recent techniques. We train the classifier and our joint-diffusion model (ensuring that both models are of similar sizes) on the CIFAR10 dataset until convergence (for 100,000 optimizer steps) [^1] Details on this experiment are provided in the Appendix. Then, we continue to optimize the models, but instead of using real data, we sample data from the diffusion model. As presented in Fig. 1, the performance of the classifier (blue line), drastically decreases as the effect of the rehearsal with generated samples. At the same time, the performance of our joint diffusion (yellow) converges after a much smaller drop in performance. Furthermore, the addition of knowledge distillation significantly decreases the degradation of both models, while keeping the trend of the joint diffusion model outperforming the classifier in terms of final performance and stability. We conclude that a combination of joint modeling and knowledge distillation is a key for knowledge retention in generative replay-base CL.
Therefore, we introduce a joint-diffusion continual learning method, that in combination with knowledge distillation and two-stage training allows for high-quality modeling in CL scenarios. In our experiments, we show that JDCL achieves state-of-the-art performance when compared with other generative replay methods, while significantly reducing the computational cost. We show that this performance gain is attributed to the higher quality of rehearsal samples when compared with the standard diffusion model. Moreover, we show that our approach can be easily applied in the continual semi-supervised setup, where thanks to the shared parametrization we can observe an alignment between labeled and unlabeled data. In this task, we show the superiority of our method over the buffer-based techniques. Finally, we also highlight the quality and stability of representations built by our method by evaluating it in a self-supervised manner. Our main contributions can be summarized as follows:
@@
  • We show that in class-incremental setup, our approach outperforms recent state-of-the-art generative replay techniques and approaches the soft-upper bound defined by a method with an infinite memory buffer, while significantly reducing the computational cost.
@@ 2 Related Work
Several works further extend the basic generative replay technique. In [68], followed by [36], authors propose to replay internal data representations instead of the original inputs. This approach improves model stability but limits its adaptation to new tasks. On the other hand, several methods drew inspiration from neuroscience and introduced techniques that mimic the human memory system by dividing the update of the model into two [24, 10] or even three stages [69].
Recently, diffusion-based generative models (DDGMs) [60, 21] have come to the forefront as a state-of-the-art generative method. Therefore, there is growing research on continual learning with DDGMs. Zajac et al. [77] provide an overview of how DDGMs work when combined with baseline CL methods. In [18] authors employ DDGMs with classifier guidance [13] as a source of rehearsal examples. This idea is further extended in [8], where authors overview different sampling strategies with DDGMs. There are several methods that explore the possibilities for efficient updates of large-scale text-to-image diffusion models. In DreamBooth [52] authors introduce a method for finetuning text-to-image DDGMs with Prior Preservation Loss, that employs generative replay to mitigate catastrophic forgetting. Similarly, [59] extend this idea with Low Rank Adaptation (LORA [22]) technique, while in [57] apply this method for object and style update. On the other hand, Basu et al. [3], show that knowledge in text-to-image models is highly localized and can be updated directly in the attention layers of the text encoder.
@@
Semi-supervised methods confront the problem of training models on large datasets where only a few samples have associated annotations [61, 76]. Recently, most of the best-performing semi-supervised methods belong to the family of hybrid methods that combine consistency regularization and pseudo-labeling  [4, 29, 61, 76, 78]. The main idea behind consistency regularization is an assumption that the model should yield similar outputs when fed perturbed versions of the same image or when its weights are slightly perturbed [31, 54, 63, 41]. On the other hand, pseudo-labeling techniques exploit the concept of using synthetic labels for unlabeled data. More precisely, this involves utilizing ”hard” labels (i.e., the arg max of the model’s output) for which the predicted class probability surpasses a predetermined threshold [74, 7, 71, 47, 32, 76]. Apart from consistency regularization and pseudo-labeling, a third group of SSL methods employs additional deep generative models  [76] such as Variational Auto-Encoder (VAE) [34, 44, 14, 39, 26], or Generative Adversarial Networks (GAN)  [55, 62, 73, 17, 37], to learn the underlying class-conditional distribution of partially labeled data.
@@ 3.2 Joint Diffusion Model
$L_{\text{class}}(\nu,\omega)=-\mathbb{E}_{\mathbf{x}_{0},y}\left[\sum_{k=0}^{K-1}\mathds{1}[y=k]\log\frac{\exp\left(\varphi_{k}\right)}{\sum_{c=0}^{K-1}\exp\left(\varphi_{c}\right)}\right]$ (6)
@@
$L_{JD}(\nu,\psi,\omega)=$ $\alpha\cdot L_{\text{class}}(\nu,\omega)-L_{0}(\nu,\psi)$
$-\sum_{t=2}^{T}L_{t,\text{diff}}(\nu,\psi)-L_{T}(\nu,\psi),$ (7)
@@ 4 Method
Figure 2: Overview of JDCL. To adapt the global model to task $\tau+1$, we first train a local joint diffusion model on task $\tau+1$ only. We then use the global model and the local model to generate a synthetic dataset consisting of samples from tasks $11,2\dots\tau+$ 1. Finally, we fine-tune the global model with the generated dataset and a combination of joint diffusion and knowledge distillation losses.
@@
One of the crucial problems in continually trained neural networks is the plasticity-stability dilemma. As presented in Fig 1, we can easily achieve high stability of our continually trained model through a self-rehearsal mechanism combined with knowledge distillation. However, to improve the plasticity, we relate to the two-stage training [10, 19] where we first fine-tune the copy of the current model using only newly available data which results in perfect plasticity. Then, we distill the adapted model with the previous one into a new global model as presented in Fig. 2.
More precisely, we first create a copy of our current joint diffusion model $p_{\nu,\psi,\omega}(x_{0:T},y)$, and train it with the currently available dataset $D_{\tau}$. We call that model local and we optimize it with the loss provided in Equation 3.2.
Following that, we move to the global stage. We begin by generating a rehearsal dataset $S_{1\dots\tau-1}$ for tasks $1$ to $\tau-1$ by sampling from the original global joint diffusion model. We then combine this dataset with a synthetic dataset $S_{\tau}$, which we generate with the local model. All the samples are generated unconditionally by the diffusion part of the joint diffusion. The labels for those generations are then assigned by the classifier part of the appropriate models.
Finally, to acquire the next global model, we fine-tune the previous global joint diffusion on the combined synthetic dataset with a joint diffusion loss $L_{JD}$ defined in Equation 3.2.
@@
$L_{JDKD}(\nu,\psi,\omega;p^{f})=$ $\alpha_{KD}\cdot L^{KD}_{\text{class}}(\nu,\omega)-L^{KD}_{0}(\nu,\psi)$
$-\sum_{t=2}^{T}L^{KD}_{t,\text{diff}}(\nu,\psi)-L^{KD}_{T}(\nu,\psi),$ (10)
@@
Finally, the training objective, which we optimize to obtain the global model for task $t$ can be expressed as:
@@
$L_{\text{ssl}}(\nu,\omega)=$ $-\mathbb{E}_{\mathbf{x}_{0},y}\{\mathds{1}[max(\varphi_{q})\geq\tau]$
$H(\hat{q_{b}},g_{\omega}(e_{\nu}(A(x_{\text{u}}))))\}$ (11)
@@
$L(\nu,\psi,\omega)=$ $\alpha\cdot L_{\text{class}}(\nu,\omega)+\alpha\cdot L_{\text{ssl}}(\nu,\omega)-L_{0}(\nu,\psi)$
$-\sum_{t=2}^{T}L_{t,\text{diff}}(\nu,\psi)-L_{T}(\nu,\psi).$
@@ 5.1 Experimental Setup
We evaluate our method on three commonly used datasets – CIFAR10, CIFAR100 [28] and ImageNet100 [12]. The first two of those datasets contain 50000 training and 10000 test images of size $32\times 32$ divided into 10 or 100 classes. ImageNet100 on the other hand contains 120000 train images in $64\times 64$ resolution from 100 different classes. For the class-incremental setup, we split the CIFAR10 and ImageNet100 datasets into 5 equal tasks and the CIFAR100 dataset into 5 or 10 equal tasks. For semi-supervised setups, we train our models on two different ratios of labeled data in the dataset i.e. 0.8% and 5%. Those ratios correspond to 4 and 25 labelled examples per class in CIFAR100, and to 40 and 250 samples for CIFAR10.
Method CIFAR-10 CIFAR-100 ImageNet100 CIFAR-10 CIFAR-100 ImageNet100
$T=5$ $T=5$ $T=10$ $T=5$ $T=5$ $T=5$ $T=10$ $T=5$
Joint 93.14 $\pm$ 0.16 72.32 $\pm$ 0.24 66.85 $\pm$ 2.25 - - - -
Continual Joint 86.41 $\pm$ 0.32 73.07 $\pm$ 0.01 64.15 $\pm$ 0.98 50.59 $\pm$ 0.35 02.90 $\pm$ 0.08 07.80 $\pm$ 0.55 06.67 $\pm$ 0.36 12.28 $\pm$ 0.07
Fine-tuning 18.95 $\pm$ 0.20 16.92 $\pm$ 0.03 9.12 $\pm$ 0.04 13.49 $\pm$ 0.18 94.65 $\pm$ 0.17 80.75 $\pm$ 0.22 87.67 $\pm$ 0.07 64.93 $\pm$ 0.00
DGR VAE 28.23 $\pm$ 3.84 19.66 $\pm$ 0.27 10.04 $\pm$ 0.17 9.54 $\pm$ 0.26 57.21 $\pm$ 9.82 42.10 $\pm$ 1.40 60.31 $\pm$ 4.80 40.46 $\pm$ 0.91
DGR+distill 27.83 $\pm$ 1.20 21.38 $\pm$ 0.61 13.94 $\pm$ 0.13 11.77 $\pm$ 0.47 43.43 $\pm$ 2.60 29.30 $\pm$ 0.40 21.15 $\pm$ 1.30 41.17 $\pm$ 0.43
RTF 30.36 $\pm$ 1.40 17.45 $\pm$ 0.28 12.80 $\pm$ 0.78 8.03 $\pm$ 0.05 51.77 $\pm$ 1.00 47.68 $\pm$ 0.80 45.21 $\pm$ 5.80 41.2 $\pm$ 0.20
MeRGAN 51.65 $\pm$ 0.40 09.65 $\pm$ 0.14 12.34 $\pm$ 0.15 - - - - -
BIR 36.41 $\pm$ 0.82 21.75 $\pm$ 0.08 15.26 $\pm$ 0.49 8.63 $\pm$ 0.19 65.28 $\pm$ 1.27 48.38 $\pm$ 0.44 53.08 $\pm$ 0.75 40.99 $\pm$ 0.36
GFR 26.70 $\pm$ 1.90 34.80 $\pm$ 0.26 21.90 $\pm$ 0.14 32.95 $\pm$ 0.35 49.29 $\pm$ 6.03 19.16 $\pm$ 0.55 17.44 $\pm$ 2.20 20.37 $\pm$ 1.47
DDGR 43.69 $\pm$ 2.60 28.11 $\pm$ 2.58 15.99 $\pm$ 1.08 25.59 $\pm$ 2.29 62.51 $\pm$ 3.84 60.62 $\pm$ 2.13 74.70 $\pm$ 1.79 49.52 $\pm$ 2.52
DGR diffusion 59.00 $\pm$ 0.57 28.25 $\pm$ 0.22 15.90 $\pm$ 1.01 23.92 $\pm$ 0.92 40.38 $\pm$ 0.32 68.70 $\pm$ 0.65 80.38 $\pm$ 1.34 54.44 $\pm$ 0.14
GUIDE 64.47 $\pm$ 0.45 41.66 $\pm$ 0.40 26.13 $\pm$ 0.29 39.07 $\pm$ 1.37 24.84 $\pm$ 0.05 44.30 $\pm$ 1.10 60.54 $\pm$ 0.82 27.60 $\pm$ 3.28
JDCL 83.69 $\pm$ 1.44 47.95 $\pm$ 0.61 29.04 $\pm$ 0.41 54.53 $\pm$ 2.15 11.13 $\pm$ 3.27 37.20 $\pm$ 0.75 59.55 $\pm$ 0.66 19.51 $\pm$ 6.27
@@
We use the same joint diffusion model architecture in all experiments on CIFAR10 and CIFAR100, and an upscaled version of it for ImageNet100 experiments to adjust to the bigger resolution of the dataset. For the UNet, we follow the implementation provided in [13]. For the classifier, we use a 2-layer feed-forward net with a hidden layer of 1024 and a LeakyRELU activation function. The detailed description of the training hyperparameters is provided in the Appendix, and in the code repository [^2] https://github.com/pskiers/Joint-Diffusion-in-Latent-Space.
@@ 5.2 Main results
To compare our method with other generative rehearsal approaches, we measure its performance on several CL benchmarks. In Tab. 1, we show that our method outperforms other evaluated generative replay techniques in terms of average accuracy after the last task. In particular, JDCL outperforms the recent state-of-the-art method–GUIDE on both average accuracy and forgetting. On CIFAR10 and ImageNet100, our method approaches the soft upper limit, calculated as a continual joint method trained with an infinitely long replay buffer. On those benchmarks, we surpass the previous state-of-the-art score by over 19 and 15 points, which is an increase of around 30% and 40% for CIFAR10 and ImageNet100 respectively.
@@
Semi-supervised continual learning
Apart from the standard supervised setup, we evaluate our method on a set of semi-supervised continual scenarios and compare it to the existing approaches. As shown in Tab. 2 JDCL outperforms other methods on all CIFAR10 setups, while also achieving comparable accuracy on CIFAR100. What’s worth noting is that JDCL is able to reach state-of-the-art performance on CIFAR10 without a memory buffer, while also in a majority of tasks sampling fewer examples from the diffusion model than the buffer size in the related works. For CIFAR100, we believe that the reduced performance of our method is caused by an extremely weak signal from the classifier, which introduces an imbalance in our joint training. To further explain the issue, we visualize how JDCL build latent representations in the SSL setup in the appendix.